OpenStudy (anonymous):

Solve. 5 = –(–z + 3) A. –8 B. 8 C. –2 D. 2

OpenStudy (fibonaccichick666):

How do you think you begin?

OpenStudy (anonymous):

\[5 = –(–z + 3)\] change the sign of both sides first

OpenStudy (micahm):

-3+2y-5z=-3 3x-y+3z=4 3x+5y-8z=6